# A Survey on RAG Meets LLMs: Towards Retrieval-Augmented Large Language Models

### v1

@@ 7. Conclusion
Retrieval-augmented generation (RAG), a cutting-edge AI technique, has achieved remarkable success across various applications, including recommendations, molecule generation, protein representation, and software engineering, owing to the powerful capabilities of retrieval in providing supplementary information to enhance generation performance. Recently, increasing efforts have been made to alleviate the limitations of large language models (LLMs), such as hallucination and out-of-date internal knowledge, by leveraging retrieval to provide the latest auxiliary information and teaching LLMs to harness the retrieved external knowledge. With the rapid advancements in retrieval-augmented large language models (RA-LLMs), there is a pressing need for a comprehensive and systematic overview. To bridge this gap, in this paper, we comprehensively review the RA-LLMs from architectures, training strategies, and applications perspectives, providing researchers with an in-depth understanding. Moreover, since the studies of RA-LLMs are still in the early stage, we also discuss the current limitations and several potential research directions for future research.

### v2

@@ 7. Conclusion
Retrieval-augmented generation (RAG), a cutting-edge AI technique, has achieved remarkable success across various applications, including recommendation, molecule generation, protein representation, and software engineering, owing to the potent capabilities of retrieval in providing supplementary information to enhance generation performance. Recently, increasing efforts have been made to alleviate the limitations of large language models (LLMs), such as hallucination and out-of-date internal knowledge, by leveraging retrieval to provide the latest auxiliary information and teaching LLMs to harness the retrieved external knowledge. With the rapid advancements in retrieval-augmented large language models (RA-LLMs), there is a pressing need for a comprehensive and systematic overview. To bridge this gap, in this paper, we comprehensively review the RA-LLMs from the perspectives of morel architecture, training strategy, and application area, providing researchers with an in-depth understanding. Moreover, since the studies of RA-LLMs are still in the early stage, we also discuss the current limitations and several potential research directions for future research.